Mr. Smith is 81 years older than his grandson Luke. In 3 years, Mr. Smith will be four times as old as Luke. How old is Luke now?
Answer by Boreal(15235)

Luke is x
Mr. Smith is x+81
x+84=4(x+3), three years later
x+84=4x+12
3x=72
x=24 years grandson
grandfather is 105
in three years, Luke will be 27 and grandfather 108.
Luke is 24 now ANSWER